A drag racer, starting from rest, speeds up for 395.0 m with an acceleration of +17.4 m/s2. A parachute then opens, slowing the car down with an acceleration of -5.18 m/s2. How fast is the racer moving 366.0 m after the parachute opens
Answer #1

Here is what I solved before, please modify the figures as per your question. Please let me know if you have further questions. Ifthis helps then kindly rate 5-stars.

A drag racer, starting from rest, speeds up for 402 m with an acceleration of +19.0 m/s2. A parachute then opens, slowing the car down with an acceleration of-7.00 m/s2. How fast is the racer moving 2.80 multiply.gif 102 m after the parachute opens?

Answer

let velocity at 402 m = v

so v= sqrt(2*a*s) = sqrt(2*19*402) = 123.596 m/s

after this a = - 7 m/s2

let final velocity = v2

so v2^2 = 123.596^2 - 2*7*280

=> v2 = 106.56 m/s